xx libjpeg

libjpeg

Federico Cipolletta, Roland Haas

2020-07-01

Abstract

This thorn provides access to the libjpeg library.

1 Using This Thorn Refer to the Cactus UserGuide, Sec. B2.2, in order to know how this thorn can be used in a compiled configuration and how to possibly linking another specific version, already installed steparately.

Note on possible ExternalLibraries’ location stripping

Each thorn contained in Cactus/arrangements/ExternalLibraries will automatically adopt the library version contained in the Cactus/arrangements/<library>/dist folder. In particular, the tarball in

Cactus/arrangements/<library>/dist is only used if either THORNxDIR is set to BUILD or is left empty and no

precompiled copy of the library is found. If another location is specified via the THORNxDIR variable in the <machine>.cfg file at compilation, then the Cactus/lib/sbin/strip-incdirs.sh script will automatically strip away (for safety reasons) the locations:

/include

/usr/include

/usr/local/include

from THORNxINCxDIRS which default to THORNxDIR/include. Therefore, if there is any need for using one already installed version of one external library, the aforementioned location should be avoided (e.g. indicating /home as the THORNxDIR will work with no problems if the required library is installed there) or should be carefully checked, in order to avoid unwanted stripping. The same stripping happens to THORNxLIBxDIRS in lib/sbin/strip-libdirs.sh with a larger list of directories:

/lib

/usr/lib

/usr/local/lib

/lib64

/usr/lib64

/usr/local/lib64

2 Parameters

3 Interfaces

General Implements: libjpeg

4 Schedule This section lists all the variables which are assigned storage by thorn ExternalLibraries/libjpeg. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE